# Non-commutative rings with infinitely many maximal subrings

### v1

@@ 4. Rings integral over their centers
Assume that $T$ has only finitely many maximal subring. Thus by Theorem 3.16, we deduce that $(1)$ holds. The first part of $(2)$ is clear by $(2)$ of Proposition 4.1. For the second part of $(2)$, let $M$ be any maximal ideal of $T$. Clearly, $T/M$ contains a copy of $K$ (and similarly $D$), thus $K$ and $D$ are absolutely algebraic fields. $(3)$ and the first part of $(4)$ are clear by Proposition 4.1. For the second part of $(4)$, note that $T/J(T)$ is a commutative ring which is integral over the field $K$. Hence we deduce that $T/J(T)$ is a zero-dimensional (reduced) ring and therefore is a von Neumann ring. Since $T$ has only finitely many maximal subrings, we infer that $T/J(T)$ has only finitely many maximal subring. Thus by [10, Proposition 2.1], we obtain that $T/J(T)$ is integral over $\mathbb{Z}_{p}$ (note, $K$ is an absolutely algebraic field). Since $T$ is algebraic over $K$, by [21, Theorem 4.20], $J(T)$ is a nil ideal of $T$. Hence if $x\in T$, then there there exist $n\in\mathbb{N}$ and $a_{0},\ldots,a_{n-1}\in\mathbb{Z}_{p}$ such that $(x+J(T))^{n}+a_{n-1}(x+J(T))^{n-1}+\cdots+a_{1}(x+J(T))+a_{0}=0$ in the ring $T/J(T)$. Thus $x^{n}+a_{n-1}x^{n-1}+\cdots+a_{1}x+a_{0}\in J(T)$. Now since $J(T)$ is nil, we deduce that there exists a natural number $m$ such that $(x^{n}+a_{n-1}x^{n-1}+\cdots+a_{1}x+a_{0})^{m}=0$, i.e., $x$ is integral over $\mathbb{Z}_{p}$. For $(5)$, assume that $T/M\cong T/N$. Thus $T/(M\cap N)\cong T/M\times T/M$. Now note that $T/M$ contains a copy of $K$, and therefore $T/M$ is an infinite field. Hence by Proposition 3.6, $T/M\times T/M$ has infinitely many maximal subrings. Hence we conclude that $T$ has infinitely many maximal subrings, which is absurd. For $(6)$, since $T/I$ contains a copy of $K$ and remains algebraic over $K$ we may assume that $I=0$. Now note that $T/J(T)$ is a reduced ring. Therefore $N(T)\subseteq J(T)$. As we see in the proof of $(4)$, $J(T)$ is nil and thus $J(T)=N(T)$. For $(7)$, note that $T/Q$ has no nonzero nil ideal, hence by $(6)$ we conclude that $J(T/Q)=0$, i.e., $Q$ is an intersection of a family of maximal (left) ideal of $T$. Thus $J(T)\subseteq Q$ and therefore by $(4)$ we immediately deduce that $T/Q$ is a field, i.e., $Q$ is a maximal (left) ideal of $T$. Finally for $(8)$, let $M$ be a maximal ideal of $T$ and $R$ be a subring of $T$. Since $(R+M)/M$ is a subring of the absolutely algebraic field $T/M$, we deduce that $(R+M)/M$ is a field. Therefore $R/(R\cap M)$ is a field. Hence $R\cap M\in Max(R)$. This immediately implies that $J(R)\subseteq J(T)\cap R$ and therefore $J(R)$ is a nil ideal (note, $J(T)$ is nil by $(6)$). ∎
@@
Let $T$ be a ring which is integral over its center. Assume that $T$ has only finitely many maximal subrings. if $P$ is a prime ideal of $T$ with $J(T/P)=0$, then either $P$ is a maximal ideal of $T$ or $T/P$ is an integral domain. In fact, either $T/P\cong\mathbb{M}_{n}(F)$ for some natural number $n>1$ and a finite field $F$ or $T/P$ is an integral domain.
@@
Since $T$ has only finitely many maximal subrings, then by $(1)$ of Proposition 3.10 we deduce that $Max_{l}^{\prime}(T)$ is finite. Hence by $J(T/P)=0$ we infer that there exist $M_{1},\ldots,M_{n}\in Max_{l}^{\prime}(T)$, $n\geq 0$ and $N=\bigcap_{i\in I}M_{i}$, where $M_{i}\in Max(R)\cap Max_{l}(T)$ such that $P=M_{1}\cap\cdots\cap M_{n}\cap N$. Thus by the comments preceding Proposition 3.10, either $P=M_{1}\cap\cdots\cap M_{n}$ or $P=N$, for $P$ is a prime ideal of $T$. If $P=M_{1}\cap\cdots\cap M_{n}$, then $T/P$ is a left $T$-submodule of the semisimple $T$-module $T/M_{1}\times\cdots\times T/M_{n}$. Therefore we conclude that $T/P$ is a simple Artinian ring. Hence $P$ is a maximal ideal of $T$. Thus assume that $P=N$. Since $T$ is integral over its center, we infer that for each $i\in I$, the division ring $T/M_{i}$ is integral over its center. Hence, if for some $i\in I$, $T/M_{i}$ is not a field, then by Theorem 3.1, we obtain that $T/M_{i}$ and therefore $T$ have infinitely many maximal subrings which is impossible. Thus for each $i\in I$, $T/M_{i}$ is a field. Clearly $T/P$ embeds in $\prod_{i\in I}T/M_{i}$ and therefore $T/P$ is a commutative (reduced prime) ring. Hence $T/P$ is an integral domain. For the final par, if $T/P$ is not an integral domain, then since $T/P$ is a simple Artinian ring, we conclude that $T/P\cong\mathbb{M}_{n}(D)$, where $D$ is a division ring and $n$ is a natural number. Now we have two cases: If $n=1$, then $T/P$ is a division ring which is integral over its center, hence by Theorem 3.1, we deduce that $T/P$ is a field, for $T/P$ has only finitely many maximal subrings. Thus $T/P$ is an integral domain which is absurd. Therefore $n>1$. We claim that $D$ is finite. For otherwise, by Theorem 3.5, we deduce that $T/P$ has infinitely many maximal subrings which is absurd. Thus $D$ is a finite field and hence we are done. ∎
@@
$(1)$ Let $P$ be a left primitive ideal of $T$. Clearly, $J(T/P)=0$ and therefore by Lemma 4.4, we conclude that either $P$ is a maximal ideal of $T$ or $T/P$ is a commutative left primitive ring. Since a commutative primitive ring is a field, we infer that $T/P$ is a field and therefore $P$ is a maximal ideal of $T$. Similarly each right primitive ideal of $T$ is a maximal ideal.
$(2)$ It is clear that each maximal ideal $P$ of $T$ is a strongly prime ideal of $T$, for $T/P$ is a simple ring and therefore has no nonzero nil ideals. Conversely, let $P$ be a strongly prime ideal of $T$, i.e., $T/P$ has no nonzero nil ideal. It is not hard to see that $P\cap C(T)$ is a prime ideal of $C(T)$ and therefore by the assumption is a maximal ideal of $C(T)$. Therefore $C(T)/(P\cap C(T))\cong(C(T)+P)/P$ is a field. Since $T$ is integral over its center, we immediately conclude that $T/P$ is integral over the field $E:=(C(T)+P)/P$. Hence by [21, Theorem 4.20], $J(T/P)$ is a nil ideal and therefore $J(T/P)=0$, for $P$ is a strongly prime ideal of $T$. Hence by Lemma 4.4, either $P$ is a maximal ideal of $T$ or $T/P$ is an integral domain. In the second case note that since $T/P$ is integral over the field $E$, we immediately conclude that $P$ is maximal too. ∎
In the next main result in this section, we see that if $T$ is a ring which is integral over its center and has only finitely many maximal subrings, then $T/J(T)$ is embeddable in $R\times S$, where $R$ is a direct product of certain absolutely algebraic fields and $S$ is a finite semisimple ring.
@@
Let $T$ be a ring which is integral over its center. Assume that $T$ has only finitely many maximal subrings. Then for each maximal ideal $M$ of $T$ either $T/M$ is an absolutely algebraic field with only finitely many maximal subrings or $T/M\cong\mathbb{M}_{n}(F)$ for some finite field $F$ and $n>1$. In particular, if $J(T)=0$, then $T$ embeds in
@@
where each $E_{i}$ is an absolutely algebraic field with only finitely many maximal subrings, $n_{j}>1$ and each $F_{j}$ is a finite field. Moreover $J$ is finite, $|I|\leq 2^{\aleph_{0}}$ and if $E_{i}$ is infinite for some $i\in I$, then for each $i^{\prime}\in I\setminus\{i\}$, $E_{i}\ncong E_{i^{\prime}}$.
@@
Let $M$ be a maximal ideal of $T$, then by assumption $T/M$ has only finitely many maximal subrings and is integral over its center. Hence by Theorem 3.1, we infer that whenever $T/M$ is a division ring, then $T/M$ is a field with only finitely many maximal subrings and therefore $T/M$ is an absolutely algebraic field, by [10, Corollary 1.5]. Hence assume that $T/M$ is not a division ring. Now by Theorem 3.5, whenever $T/M$ is an infinite ring, then $T/M$ has infinitely many maximal subrings which is impossible. Thus $T/M$ is a finite ring. Therefore $T/M\cong\mathbb{M}_{n}(F)$, where $F$ is a finite ring and $n$ is a natural number. Since $T/M$ is not a division ring in this case, we deduce that $n>1$. Also note that since $T$ has only finitely many maximal subrings, then $T$ has only finitely many distinct maximal ideals $M$ for them the ring $T/M$ is not a field, by $(5)$ of Proposition 3.10, say $M_{1},\ldots,M_{k}$ (and assume that $T/M_{i}\cong\mathbb{M}_{n_{i}}(F_{i})$, where $F_{i}$ is a finite field and $n_{i}>1$). For the next part, by $(8)$ of Proposition 3.10, we have $0=J(T)=\bigcap_{M\in Max(T)}M$. Hence we conclude that $0=J(T)=M_{1}\cap\cdots\cap M_{k}\cap\bigcap_{i\in I}N_{i}$, where for each $i\in I$, $T/N_{i}$ is an absolutely algebraic field with only finitely many maximal subrings, by the first part of the theorem. Clearly, we may assume that for $i\neq j$ in $I$, $N_{i}\neq N_{j}$. Thus $T$ embeds in $T/M_{1}\times\cdots\times T/M_{k}\times\prod_{i\in I}T/N_{i}$. Clearly, by Proposition 4.2, we have $|Max(T)|\leq 2^{\aleph_{0}}$ and therefore $|I|\leq 2^{\aleph_{0}}$. Finally we show that if $i\neq j$ in $I$ and $T/N_{i}$ is infinite, then $T/N_{i}\ncong T/N_{j}$. Note that if $T/N_{i}\cong T/N_{j}$, then by Proposition 3.6, we deduce that $T/(N_{i}\cap N_{j})\cong T/N_{i}\times T/N_{i}$ has infinitely many maximal subrings which is impossible. Hence we are done (note that clearly $|J|=k$ is finite). ∎
In the next section we exactly determine when the ring $T=\prod_{i\in I}\mathbb{M}_{n_{i}}(E_{i})$, where each $E_{i}$ is a field and $n_{i}\in\mathbb{N}$, has only finitely many maximal subrings. In the following we obtain some results for Artinian/Noetherian rings. First we need the following remark.
@@
It is not hard to see that if $S$ is a commutative ring with nonzero characteristic, say $n$, $Max(S)=\{M_{1},\ldots,M_{k}\}$ and $J(S)$ is a nil ideal of $S$ (in particular, if $S$ is a commutative Artinian ring with nonzero characteristic $n$), then $S$ is integral over $\mathbb{Z}_{n}$ if and only if each $S/M_{i}$ is an absolutely algebraic field.
@@
$(1)$ First note that by [8, Propositions 1.4 and 2.4] and the previous remark, $C$ has a maximal ideal $M$ such that $C/M$ is a field which is either uncountable or of zero characteristic and or is not absolutely algebraic field, respectively. Hence in any cases, $K:=C/M$ is not an absolutely algebraic field. By [11, Proposition 1.2], $T$ has a prime ideal $Q$ such that $Q\cap C=M$. Clearly, $K=C/M\subseteq C(T/Q)$ and $T/Q$ is integral over $K$ and therefore $T/Q$ is integral over its center. Thus by $(2)$ of Theorem 4.3, we conclude that $T/Q$ has infinitely many maximal subrings. Hence $T$ has infinitely many maximal subrings.
$(2)$ First assume that $C$ is an integral domain. By the proof of [9, Corollary 2.7], for each maximal ideal $M$ of $C$, there exists a natural number $n$, such that $C/M^{n}$ is an uncountable ring. Since $C/M^{n}$ is a zero dimensional noetherian local ring with only maximal ideal $M/M^{n}$, by the first part of the proof of $(1)$, we deduce that $C/M$ is an uncountable field and similar to the proof of $(1)$, we conclude that $T$ has infinitely many maximal subrings. If $C$ is not an integral domain, then by the proof of [9, Theorem 2.9], $C$ has a minimal prime ideal $P$ such that $|C/P|=|C|$. By [11, Proposition 1.2], $T$ has a prime ideal $Q$ such that $Q\cap C=P$. Clearly $T/Q$ is integral over $C/P\subseteq C(T/Q)$ (and therefore $T/Q$ is integral over its center). Similar to the first part of the proof of $(2)$, for each maximal ideal $M/P$ of $C/P$, the field $(C/P)/(M/P)$ is uncountable. Since $T/Q$ is integral over $C/P$, by [11, Proposition 1.2], $T/Q$ has a prime ideal $N/Q$ such that $(N/Q)\cap(C/P)=M/P$ and clearly $(T/Q)/(N/Q)$ is integral over $K:=(C/P)/(M/P)\subseteq C((T/Q)/(N/Q))$ (and therefore $(T/Q)/(N/Q)$ is integral over its center). Now since $(T/Q)/(N/Q)$ contains an uncountable field $K=(C/P)/(M/P)$, we immediately conclude that $(T/Q)/(N/Q)$ has infinitely many maximal subrings by $(2)$ of Theorem 4.3. Thus $T$ has infinitely many maximal subrings too. ∎
@@
Assume that $T$ has only finitely many maximal subrings, then by $(4)$ of Theorem 4.3, we infer that $T/J(T)$ is a von Neumann regular commutative ring and $T$ is integral over $\mathbb{Z}_{p}$, where $p=Char(K)$ is a prime number. Now note that there exists a natural number $m$ such that $J(T)^{m}=0$, for $T$ is a left Artinian ring. Since $T/J(T)$ is an Artinian commutative ring with only finitely many maximal subrings, by $(2)$ of [10, Theorem 2.5], we deduce that $T/J(T)$ is a countable ring. Hence by Lemma 2.8, $T$ is a countable ring. ∎
@@
First note that since $T$ is an algebraic $K$-algebra over a field $K$, we infer that $J(T)$ is a nil ideal of $T$, by [21, Theorem 4.20]. Therefore by [21, Theorem 10.30], $J(T)$ is a nilpotent ideal of $T$, for $T$ is a left Noetherian ring. Hence we conclude that $T$ is a left Artinian ring by [21, Theorem 4.15]. Thus by Theorem 4.9, we deduce the assertion. ∎

### v2

@@ 4. Rings integral over their centers
Assume that $T$ has only finitely many maximal subring. Thus by Theorem 3.14, we deduce that $(1)$ holds. The first part of $(2)$ is clear by Proposition 4.1(2)(2). For the second part of $(2)$, let $M$ be any maximal ideal of $T$. Clearly, $T/M$ contains a copy of $K$ (and similarly $D$), thus $K$ and $D$ are absolutely algebraic fields. Consequently, $\aleph_{0}\leq|K|\leq|T/M|\leq\aleph_{0}$ and thus the equalities hold. $(3)$ and the first part of $(4)$ are evident by Proposition 4.1. For the second part of $(4)$, note that $T/J(T)$ is a commutative ring which is integral over the field $K$. Hence we deduce that $T/J(T)$ is a zero-dimensional (reduced) ring and a fortiori, it is a von Neumann ring. Since $T$ has only finitely many maximal subrings, we infer that $T/J(T)$ has only finitely many maximal subring. Thus by [10, Proposition 2.1], we obtain that $T/J(T)$ is integral over $\mathbb{Z}_{p}$ (note, $K$ is an absolutely algebraic field). Since $T$ is algebraic over $K$, by [21, Theorem 4.20], $J(T)$ is a nil ideal of $T$. Now let $x\in T$, then there exist $n\in\mathbb{N}$ and $a_{0},\ldots,a_{n-1}\in\mathbb{Z}_{p}$ such that $(x+J(T))^{n}+a_{n-1}(x+J(T))^{n-1}+\cdots+a_{1}(x+J(T))+a_{0}=0$ in the ring $T/J(T)$. Thus $x^{n}+a_{n-1}x^{n-1}+\cdots+a_{1}x+a_{0}\in J(T)$. From the fact that $J(T)$ is nil, we deduce that there exists a natural number $m$ such that $(x^{n}+a_{n-1}x^{n-1}+\cdots+a_{1}x+a_{0})^{m}=0$, i.e., $x$ is integral over $\mathbb{Z}_{p}$. For $(5)$, assume that $T/M\cong T/N$. Thus $T/(M\cap N)\cong T/M\times T/M$. Now note that $T/M$ contains a copy of $K$, and therefore $T/M$ is an infinite field. Hence by Proposition 3.6, $T/M\times T/M$ has infinitely many maximal subrings. It follows that $T$ has infinitely many maximal subrings, which is absurd. For $(6)$, nothing that $T/I$ contains a copy of $K$ and remains algebraic over $K$, thus we may assume that $I=0$. Because $T/J(T)$ is a reduced ring, it is obvious that $N(T)\subseteq J(T)$. Consequently, as we see in the proof of $(4)$, $J(T)$ is nil and then $J(T)=N(T)$. For $(7)$, note that $T/Q$ has no nonzero nil ideal, hence by $(6)$ we conclude that $J(T/Q)=0$, i.e., $Q$ is an intersection of a family of maximal (left) ideal of $T$. Thus $J(T)\subseteq Q$ and then by $(4)$, we deduce that $T/Q$ is a field, i.e., $Q$ is a maximal (left) ideal of $T$. Finally for $(8)$, let $M$ be a maximal ideal of $T$ and $R$ be a subring of $T$. Since $(R+M)/M$ is a subring of the absolutely algebraic field $T/M$, we obtain that $(R+M)/M$ is a field. It follows that $R/(R\cap M)$ is a field, i.e., $R\cap M\in\operatorname{Max}(R)$. This implies that $J(R)\subseteq J(T)\cap R$ and at once $J(R)$ is a nil ideal (note, $J(T)$ is nil by $(6)$). ∎
@@
Let $T$ be a ring which is integral over its center. Assume that $T$ has only finitely many maximal subrings. If $P$ is a prime ideal of $T$ with $J(T/P)=0$, then either $P$ is a maximal ideal of $T$ or $T/P$ is an integral domain. In fact, either $T/P\cong\mathbb{M}_{n}(F)$ for some natural number $n>1$ and finite field $F$ or $T/P$ is an integral domain.
@@
Since $T$ has only finitely many maximal subrings, then by Proposition 3.10(1)(1) we deduce that $\operatorname{{Max}_{\ell}}^{\prime}(T)$ is finite. Hence by $J(T/P)=0$ we infer that there exist $M_{1},\ldots,M_{n}\in\operatorname{{Max}_{\ell}}^{\prime}(T)$, $n\geq 0$ and $N=\bigcap_{i\in I}M_{i}$, where $M_{i}\in\operatorname{Max}(R)\cap\operatorname{{Max}_{\ell}}(T)$ such that $P=M_{1}\cap\cdots\cap M_{n}\cap N$. Thus by the comments preceding Proposition 3.10, either $P=M_{1}\cap\cdots\cap M_{n}$ or $P=N$, because $P$ is a prime ideal of $T$. If $P=M_{1}\cap\cdots\cap M_{n}$, then $T/P$ is a left $T$-submodule of the semisimple $T$-module $T/M_{1}\times\cdots\times T/M_{n}$. Consequently $T/P$ is a simple Artinian ring. Hence $P$ is a maximal ideal of $T$. Thus assume that $P=N$. Since $T$ is integral over its center, we infer that for each $i\in I$, the division ring $T/M_{i}$ is integral over its center. Hence, if for some $i\in I$, $T/M_{i}$ is not a field, then by Proposition 3.1, we obtain that $T/M_{i}$ and therefore $T$ have infinitely many maximal subrings which is impossible. Thus for each $i\in I$, $T/M_{i}$ is a field. Clearly $T/P$ embeds in $\prod_{i\in I}T/M_{i}$ and therefore $T/P$ is a commutative (reduced prime) ring. It follows that $T/P$ is an integral domain, because $P$ is prime. For the last part, if $T/P$ is not an integral domain, then since $T/P$ is a simple Artinian ring, we infer that $T/P\cong\mathbb{M}_{n}(D)$, where $D$ is a division ring and $n$ is a natural number. Now we have two cases: If $n=1$, then $T/P$ is a division ring which is integral over its center, hence by Proposition 3.1, we deduce that $T/P$ is a field, for $T/P$ has only finitely many maximal subrings. Thus $T/P$ is an integral domain which is absurd. Now suppose that $n>1$, we claim that $D$ is finite. Otherwise, by Theorem 3.5, we deduce that $T/P$ has infinitely many maximal subrings which is absurd. Thus $D$ is a finite field, as desired. ∎
@@
$(1)$ Let $P$ be a left primitive ideal of $T$. Clearly, $J(T/P)=0$ and therefore by Lemma 4.4, we conclude that either $P$ is a maximal ideal of $T$ or $T/P$ is a commutative left primitive ring. Since a commutative primitive ring is a field, we infer that $T/P$ is a field and consequently $P$ is a maximal ideal of $T$. Similarly each right primitive ideal of $T$ is a maximal ideal.
$(2)$ It is clear that each maximal ideal $P$ of $T$ is a strongly prime ideal of $T$, for $T/P$ is a simple ring and therefore has no nonzero nil ideals. Conversely, let $P$ be a strongly prime ideal of $T$, i.e., $T/P$ has no nonzero nil ideal. It is not hard to see that $P\cap C(T)$ is a prime ideal of $C(T)$. It follows that by the assumption $P\cap C(T)$ is a maximal ideal of $C(T)$. This implies that $C(T)/(P\cap C(T))\cong(C(T)+P)/P$ is a field. Since $T$ is integral over its center, we conclude that $T/P$ is integral over the field $E:=(C(T)+P)/P$. Hence by [21, Theorem 4.20], $J(T/P)$ is a nil ideal and therefore $J(T/P)=0$, because $P$ is a strongly prime ideal of $T$. Consequently by Lemma 4.4, either $P$ is a maximal ideal of $T$ or $T/P$ is an integral domain. In the second case note that since $T/P$ is integral over the field $E$, we obtain that $P$ is maximal. ∎
The next main result in this section shows that if $T$ is a ring that is integral over its center and has only finitely many maximal subrings, then $T/J(T)$ is embeds into a ring of the form $R\times S$, where $R$ is a direct product of certain absolutely algebraic fields and $S$ is a finite semisimple ring.
@@
Let $T$ be a ring which is integral over its center. Assume that $T$ has only finitely many maximal subrings. Then for each maximal ideal $M$ of $T$ either $T/M$ is an absolutely algebraic field with only finitely many maximal subrings or $T/M\cong\mathbb{M}_{n}(F)$ for some finite field $F$ and $n>1$. In particular, if $J(T)=0$, then $T$ embeds into
@@
where each $E_{i}$ is an absolutely algebraic field with only finitely many maximal subrings, $n_{j}>1$ and each $F_{j}$ is a finite field. Moreover $J$ is finite, $|I|\leq 2^{\aleph_{0}}$ and if $E_{i}$ is infinite for some $i\in I$, then for each $i^{\prime}\in I\setminus\{i\}$, $E_{i}\ncong E_{i^{\prime}}$, as rings.
@@
Let $M$ be a maximal ideal of $T$, then by the assumption $T/M$ has only finitely many maximal subrings and is integral over its center. Hence by Theorem 3.1, we infer that whenever $T/M$ is a division ring, then $T/M$ is a field with only finitely many maximal subrings and therefore $T/M$ is an absolutely algebraic field, by [10, Corollary 1.5]. Thus assume that $T/M$ is not a division ring. Now by Theorem 3.5, whenever $T/M$ is an infinite ring, then $T/M$ has infinitely many maximal subrings which is impossible. Thus $T/M$ is a finite ring. Therefore $T/M\cong\mathbb{M}_{n}(F)$, where $F$ is a finite ring and $n$ is a natural number. Since $T/M$ is not a division ring in this case, we deduce that $n>1$. Also note that since $T$ has only finitely many maximal subrings, then $T$ has only finitely many distinct maximal ideals $M$ for them the ring $T/M$ is not a field, by Proposition 3.10(5)(5), say $M_{1},\ldots,M_{k}$ (and suppose that $T/M_{i}\cong\mathbb{M}_{n_{i}}(F_{i})$, for some finite field $F_{i}$ and $n_{i}>1$). For the next part, by Proposition 3.10(8)(8), we have $0=J(T)=\bigcap_{M\in\operatorname{Max}(T)}M$. Consequently by the first part of the proof of the theorem, $0=J(T)=M_{1}\cap\cdots\cap M_{k}\cap\bigcap_{i\in I}N_{i}$, where for each $i\in I$, $E_{i}:=T/N_{i}$ is an absolutely algebraic field with only finitely many maximal subrings. Clearly, we may assume that for $i\neq j$ in $I$, $N_{i}\neq N_{j}$. Thus $T$ embeds in $T/M_{1}\times\cdots\times T/M_{k}\times\prod_{i\in I}T/N_{i}$. It follows that from Proposition 4.2, $|\operatorname{Max}(T)|\leq 2^{\aleph_{0}}$ and then $|I|\leq 2^{\aleph_{0}}$. Finally we show that if $i\neq j$ in $I$ and $T/N_{i}$ is infinite, then $T/N_{i}\ncong T/N_{j}$. Note that if $T/N_{i}\cong T/N_{j}$, then by Proposition 3.6, we deduce that $T/(N_{i}\cap N_{j})\cong T/N_{i}\times T/N_{i}$ has infinitely many maximal subrings which is impossible. Hence we are done (note that clearly $|J|=k$ is finite). ∎
In the next section we precisely determine when a ring of the form $T=\prod_{i\in I}\mathbb{M}_{n_{i}}(E_{i})$, where each $E_{i}$ is a field and $n_{i}\in\mathbb{N}$, has only finitely many maximal subrings. Below we derive some related results for Artinian/Noetherian rings. First we require the following remark.
@@
It is not hard to see that if $S$ is a commutative ring with nonzero characteristic, say $n$, $\operatorname{Max}(S)=\{M_{1},\ldots,M_{k}\}$ and $J(S)$ is a nil ideal of $S$ (in particular, if $S$ is a commutative Artinian ring with nonzero characteristic $n$), then $S$ is integral over $\mathbb{Z}_{n}$ if and only if each $S/M_{i}$ is an absolutely algebraic field.
@@
$(1)$ First note that by [8, Propositions 1.4 and 2.4] and the previous remark, $C$ has a maximal ideal $M$ such that $C/M$ is a field which is either uncountable or of zero characteristic and or is not absolutely algebraic field, respectively. Hence in any cases, $K:=C/M$ is not an absolutely algebraic field. By [11, Proposition 1.2], $T$ has a prime ideal $Q$ such that $Q\cap C=M$. Clearly, $K=C/M\subseteq C(T/Q)$ and $T/Q$ is integral over $K$ and therefore $T/Q$ is integral over its center. Thus by Theorem 4.3(2)(2), we conclude that $T/Q$ has infinitely many maximal subrings. Hence $T$ has infinitely many maximal subrings.
$(2)$ First assume that $C$ is an integral domain. By the proof of [9, Corollary 2.7], for each maximal ideal $M$ of $C$, there exists a natural number $n$, such that $C/M^{n}$ is an uncountable ring. Since $C/M^{n}$ is a zero dimensional noetherian local ring with only maximal ideal $M/M^{n}$, by the first part of the proof of $(1)$, we deduce that $C/M$ is an uncountable field and similar to the proof of $(1)$, we conclude that $T$ has infinitely many maximal subrings. If $C$ is not an integral domain, then by the proof of [9, Theorem 2.9], $C$ has a minimal prime ideal $P$ such that $|C/P|=|C|$. By [11, Proposition 1.2], $T$ has a prime ideal $Q$ such that $Q\cap C=P$. Clearly $T/Q$ is integral over $C/P\subseteq C(T/Q)$ (and therefore $T/Q$ is integral over its center). Similar to the first part of the proof of $(2)$, for each maximal ideal $M/P$ of $C/P$, the field $(C/P)/(M/P)$ is uncountable. Since $T/Q$ is integral over $C/P$, by [11, Proposition 1.2], $T/Q$ has a prime ideal $N/Q$ such that $(N/Q)\cap(C/P)=M/P$ and clearly $(T/Q)/(N/Q)$ is integral over $K:=(C/P)/(M/P)\subseteq C((T/Q)/(N/Q))$ (and therefore $(T/Q)/(N/Q)$ is integral over its center). Now since $(T/Q)/(N/Q)$ contains an uncountable field $K=(C/P)/(M/P)$, we immediately conclude that $(T/Q)/(N/Q)$ has infinitely many maximal subrings by Theorem 4.3(2)(2). Thus $T$ has infinitely many maximal subrings too. ∎
@@
Assume that $T$ has only finitely many maximal subrings, then by Theorem 4.3(4)(4), we infer that $T/J(T)$ is a von Neumann regular commutative ring and $T$ is integral over $\mathbb{Z}_{p}$, where $p=\operatorname{Char}(K)$ is a prime number. Now note that there exists a natural number $m$ such that $J(T)^{m}=0$, for $T$ is a left Artinian ring. Since $T/J(T)$ is an Artinian commutative ring with only finitely many maximal subrings, by [10, Theorem 2.5(2)(2)], we deduce that $T/J(T)$ is a countable ring. Hence by Lemma 2.8, $T$ is a countable ring. ∎
@@
First note that since $T$ is an algebraic $K$-algebra over a field $K$, we infer that $J(T)$ is a nil ideal of $T$, by [21, Theorem 4.20]. Therefore by [21, Theorem 10.30], $J(T)$ is a nilpotent ideal of $T$, because $T$ is a left Noetherian ring. Hence $T$ is a left Artinian ring, by [21, Theorem 4.15]. Thus by Theorem 4.9, we deduce the assertion. ∎